OpenStudy (anonymous):

A 2000 kg truck is driving at 38 m/s along a highway when a parachute of mass 96 kg falls vertically and lands on the bed of the truck. Assuming the truck driver does not accelerate, what is the velocity of the truck after the parachute lands? ____m/s

OpenStudy (anonymous):

okay I got \[v_f=36.2595 m/s\] \[P_B=mv=2000kg(38m/s)=76000kg m/s\] \[P_A=76000=2096(v_f)\] \[v_f=36.2595m/s\]

OpenStudy (theeric):

Looks good! :)

OpenStudy (anonymous):

it was right!

OpenStudy (theeric):

Awesome! :)